Calculate the mass percent of solute in each solution. 


Calculate the mass percent of 2.77g KCl dissolved in 46.1 g H₂O. 

Calculate the mass percent of 25.9 g KNO₃ dissolved in 823 g H₂O. 
Answer #1

1) 2.77g KCl + 46.1g H2O = 48.87

    2.77/48.87 (100) = 5.67%


2) 25.9gKNO3 + 823gH2O = 848.9

    25.9/823 (100) = 3.05%

answered by: Chloë Dick